A particle of mass 'm' carry a charge of '-q1' starts moving around a fixed charge q2 along a circular path of radius r. What will be the period of revolution? 
 

Grade:12

2 Answers

Arun
5 years ago
Dear student
 
As we know
 

Fe=Centripital force

k*q1*q2/r2 = Mrw2  [w=omega]

w2= k*q1*q2/Mr3

(2Π/T )2= k*q1*q2/Mr3

T2 = 4ΠMr3/k*q1*q2

T = (4ΠMr3/k*q1*q2)1/2

In case of any difficulty, please feel free to contact.

 

Regards

Arun (askIITians forum expert)

Khimraj
5 years ago
Kq1*q2/r2 = mv2/r
v = \sqrt{K*q1*q2/m*r}
so T = 2\pir/v = 2\pir(\sqrt{m*r/K*q1*q2}
Hope it clears.